OpenStudy (anonymous):

I have to solve the equation "log(3x^2-3x-35)=0". I know that the answer is "x=-3 v x=4", but I'd like to know how to figure that out.

OpenStudy (campbell_st):

well raise each side of the equation as a power of the base of the log... use the rule \[e^{\ln(x)} = x\] so I'll assume base e logs \[e^{\ln(3x^2 -3x -35)} = e^0\] which becomes \[3x^3 -3x -35 = 1\] hope you can solve the quadratic from here